If 8.66 X 1021 atoms are removed from 0.055 mole of iron, what mass of iron remains?
1
Expert's answer
2020-05-01T14:40:48-0400

Solution.

"\Delta N = 8.66 \sdot 10^{21} atoms;"

"\nu_0 = 0.055 mol;"

We make a proportion:

"1 mol ---- 6.02\sdot10^{23} atoms;"

"0.055 mol ---x atoms;"

"x =\dfrac{0.055mol\sdot6.02\sdot10^{23}}{1 mol} =33.11\sdot10^{21} atoms;"

"N = 33.11\sdot10^{21} - 8.66\sdot10^{21} = 24.45\sdot10^{21} atoms;"

Let's find how many moles of iron are left:

"1 mol ---- 6.02\sdot10^{23} atoms;"

"\nu mol----24.45\sdot10^{21} atoms;"

"\nu = \dfrac{24.45\sdot10^{21}atoms \sdot1mol}{6.02\sdot10^{23} atoms} = 0.041mol;"

"\nu =\dfrac{m}{M}; \implies m=\nu\sdot M;"

"m = 0.041mol\sdot56g/mol =2.27g;"

Answer: "m =2.27g;"